OpenStudy (anonymous):

Could someone help me learn how to do this?? I will give a medal and fan to whoever help me learn the best! Which is the simplified form of 16a^8/12a^2?

OpenStudy (anonymous):

@robz8 ??

OpenStudy (anonymous):

eazy peazy what is a common factor in 16 and 12?

OpenStudy (anonymous):

4!

OpenStudy (anonymous):

@robz8

OpenStudy (anonymous):

so divide the numerator and denominator by 4 to simplify the coefficients

OpenStudy (anonymous):

and with variables with exponents, just subtract the exponent in the denominator from the exponent in the numerator

OpenStudy (jhannybean):

\[\large \frac{16a^8}{12a^2}= \frac{16}{12}*\frac{a^8}{a^2}\] So you said 4 goes into 16 and 12... you have \[\large \frac{4}{3}*a^{8-2}\]

OpenStudy (anonymous):

^what she said

OpenStudy (anonymous):

Thank you both!!:)

OpenStudy (jhannybean):

no problem :) rob had it.